OpenStudy (jewelliasmith1):

your school is planning a fundraising dinner. the expense for this event must not exceed $2,475.00. The team organizing the event has calculated that the cost per adult guest will be $18.00 and the cost per child will be $9.00. The venue can hold no more than 150 guest. write a system of inequalities that represent this situation?

OpenStudy (jewelliasmith1):

@skullpatrol

OpenStudy (jewelliasmith1):

no i don't understand what i'm supposed to be doing

OpenStudy (perfecta):

x is the number of adults attending the event y is the number of children attending the event. x + y <= 150 is one equation. 18x + 9y <= 2475 is the other equation. x >= 0 is another equation that is implied because the number of adults can't be less than 0. y >= 0 is another equation for the same reason as in x >= 0.

OpenStudy (adonine):

x is the number of adults attending the event y is the number of children attending the event. x + y <= 150 is one equation. 18x + 9y <= 2475 is the other equation. x >= 0 is another equation that is implied because the number of adults can't be less than 0. y >= 0 is another equation for the same reason as in x >= 0.

OpenStudy (jewelliasmith1):

so what equation do i use to salve the question

OpenStudy (tylerocks):

what grade is it

OpenStudy (jewelliasmith1):

10th

OpenStudy (tylerocks):

ok

OpenStudy (jewelliasmith1):

do you think you can help me

OpenStudy (tylerocks):

a little

OpenStudy (tylerocks):

x/y

OpenStudy (tylerocks):

thats what u have to do

OpenStudy (tylerocks):

who likes history

OpenStudy (jewelliasmith1):

so x= adults and y= children

OpenStudy (jewelliasmith1):

@TYLEROCKS

OpenStudy (tylerocks):

yes

OpenStudy (jewelliasmith1):

so can u tell me how i write the equation